# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v1

@@ Abstract
Model-based Offline Reinforcement Learning trains policies based on offline datasets and model dynamics, without direct real-world environment interactions. However, this method is inherently challenged by distribution shift. Previous approaches have primarily focused on tackling this issue directly leveraging off-policy mechanisms and heuristic uncertainty in model dynamics, but they resulted in inconsistent objectives and lacked a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two key components: model bias and policy shift. We provide both theoretical insights and empirical evidence to demonstrate how these factors lead to inaccuracies in value function estimation and impose implicit restrictions on policy learning. To address these challenges, we derive adjustment terms for model bias and policy shift within a unified probabilistic inference framework. These adjustments are seamlessly integrated into the vanilla reward function to create a novel Shifts-aware Reward (SAR), aiming at refining value learning and facilitating policy training. Furthermore, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate the SAR for policy optimization. Empirically, we show that SAR effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks, underscoring its practical effectiveness and validating our theoretical analysis.
footnotetext: footnotetext: Preprint. Under Review.
Introduction
@@
Offline reinforcement learning (RL) (Lange, Gabel, and Riedmiller 2012; Levine et al. 2020) learns policies from a pre-collected dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach showcases remarkable potential in data-driven decision scenarios (Emerson, Guy, and McConville 2023; Sinha, Mandlekar, and Garg 2022) where the exploration is costly or hazardous. The model-based framework (Luo et al. 2018) is particularly effective for offline RL (Yu et al. 2020, 2021; Sun et al. 2023; Li et al. 2024). It involves learning models of the environment from the given dataset and utilizing these models to generate data for policy training. Nevertheless, directly learning a policy using offline and model-generated data introduces the distribution shift challenge. This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
@@
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO (Yu et al. 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE (Sun et al. 2023) penalizes value learning of model data according to the uncertainty of off-policy methods in online RL and the Bellman Q-function estimated by ensemble models. RAMBO (Rigter, Lacerda, and Hawes 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions. However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
@@
Figure 1: The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) to the adjacent grid, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift. We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima. Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence. Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data. Furthermore, it learns a transition classifier and an action classifier to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of value function, while policy shift imposes an implicit policy limitation to hinder policy convergence.
  • We upgrade the vanilla reward into a shifts-aware reward with model bias adjustment and policy shift modification. This shifts-aware reward is used for policy learning to mitigate distribution shift in model-based offline RL.
@@
  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks.
Related Work
@@
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin, Yang, and Wang 2021). Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift. For instance, BEAR (Kumar et al. 2019) and BRAC (Wu, Tucker, and Nachum 2019) constrain the learned policy by minimizing different divergences. BCQ (Fujimoto, Meger, and Precup 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions. CQL (Kumar et al. 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC (Fujimoto and Gu 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto, Hoof, and Meger 2018) objective. Similarly, EDAC (An et al. 2021) and PBRL (Bai et al. 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks. MCQ (Lyu et al. 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
@@
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al. 2021), achieving high data efficiency (Li et al. 2024). MOPO (Yu et al. 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL (Kidambi et al. 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO (Yu et al. 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. RAMBO (Rigter, Lacerda, and Hawes 2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP (Jeong et al. 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al. 2018). Lastly, MOBILE (Sun et al. 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models.
Preliminaries
MDP.
@@
We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
MBPO.
@@
Similar to previous works (Yu et al. 2020, 2021; Rigter, Lacerda, and Hawes 2022; Sun et al. 2023), we focus on model-based policy optimization (MBPO) (Janner et al. 2019), which employs an actor-critic RL algorithm. MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$. Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
Trajectory Formulation.
@@
To demystify the distribution shift, We consider the trajectory formulation of optimization objectives. Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$. This is given by:
@@
Shifts-aware Policy Optimization
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shift.
Demystifying Distribution Shift
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Furthermore, we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence. The supplementary analysis is in Appendix A1.
Shift Weighting.
The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
@@
Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following equation:
@@
Negative Impact of Shift Weighting.
We analyze the impact of shift weighting from model bias and policy shift.
(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from actual environment dynamics $p$. In this context, the training objective $\mathcal{J}(\pi)$ is:
@@
Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $q$ overestimates or underestimates the likelihood of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. This essentially introduces a bias for the optimization objective, resulting in wrong training directions.
@@
Toy Examples.
To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
These analyses and empirical evidence highlight the critical role of addressing model bias and policy shift in achieving accurate value learning and efficient policy training.
Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We further examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
Shifts-aware Reward.
We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
@@
Modal Bias Adjustment.
We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
@@
Policy Shift Modification.
To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$. Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
@@
Shifts-aware Model-based Offline RL

### v2

@@ Abstract
Model-based offline reinforcement learning trains policies using pre-collected datasets and learned environment models, eliminating the need for direct real-world environment interaction. However, this paradigm is inherently challenged by distribution shift (DS). Existing methods address this issue by leveraging off-policy mechanisms and estimating model uncertainty, but they often result in inconsistent objectives and lack a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two fundamental components: model bias and policy shift. Our theoretical and empirical investigations reveal how these factors distort value estimation and restrict policy optimization. To tackle these challenges, we derive a novel Shifts-aware Reward (SAR) through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training. Building on this, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate SAR for policy optimization. Empirical experiments show that SAR effectively mitigates DS, and SAMBO-RL achieves superior or comparable performance across various benchmarks, underscoring its effectiveness and validating our theoretical analysis.
Machine Learning, ICML
1 Introduction
@@
Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach shows remarkable potential in data-driven decision scenarios (Emerson et al., 2023; Sinha et al., 2022) where exploration is costly or hazardous. The model-based framework (Luo et al., 2018) is particularly effective for offline RL (Yu et al., 2020, 2021; Sun et al., 2023; Li et al., 2024). It involves learning models of the environment from the offline dataset and utilizing these models to generate data for policy training. However, directly learning a policy using offline and synthetic data introduces the distribution shift challenge. This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
@@
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO (Yu et al., 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE (Sun et al., 2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models. RAMBO (Rigter et al., 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions. However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
Figure 1: The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells $S-1$ or $S+1$, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives undiscounted rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift. We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
@@
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima. Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence. Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data. Furthermore, it learns a transition classifier and action classifiers to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of the value function, while policy shift imposes an implicit policy limitation that hinders policy convergence.
  • We upgrade the vanilla reward into a shifts-aware reward incorporating model bias adjustment and policy shift modification. This modified reward is used for policy learning to mitigate distribution shift in model-based offline RL.
@@
  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior or comparable performance across various benchmarks.
2 Related Work
@@
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin et al., 2021). Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift. For instance, BEAR (Kumar et al., 2019) and BRAC (Wu et al., 2019) constrain the learned policy by minimizing different divergences. BCQ (Fujimoto et al., 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions. CQL (Kumar et al., 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC (Fujimoto & Gu, 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto et al., 2018) objective. Similarly, EDAC (An et al., 2021) and PBRL (Bai et al., 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks. MCQ (Lyu et al., 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
@@
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al., 2021), achieving high data efficiency (Li et al., 2024). MOPO (Yu et al., 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL (Kidambi et al., 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO (Yu et al., 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. TT (Janner et al., 2021) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter et al., 2022) adopts an adversarial approach, training policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP (Jeong et al., 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al., 2018). Lastly, MOBILE (Sun et al., 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble learned dynamic models.
3 Preliminaries
@@
MDP. We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
@@
MBPO. Similar to previous works (Yu et al., 2020, 2021; Rigter et al., 2022; Sun et al., 2023), we focus on model-based policy optimization (MBPO) (Janner et al., 2019), which employs an actor-critic RL algorithm. MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$. Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
@@
Trajectory Formulation. To demystify the distribution shift, We consider the trajectory formulation of optimization objectives. Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$. This is given by:
@@
4 Shifts-aware Policy Optimization
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shifts.
4.1 Demystifying Distribution Shift
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Then we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence. The supplementary analysis is in Appendix A.1.
Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
@@ 4.1 Demystifying Distribution Shift
Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following:
@@
Negative Impact of Shift Weighting. We analyze the impact of shift weighting from model bias and policy shift.
(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from the actual environment dynamics $p$. In this context, the training objective $\mathcal{J}(\pi)$ is:
@@
Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $q$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. This essentially introduces bias for the optimization objective, resulting in wrong training directions.
@@
Toy Examples. To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
4.2 Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
Shifts-aware Reward. We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
@@ 4.2 Shifts-aware Reward via Probabilistic Inference
Modal Bias Adjustment. We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
@@
Policy Shift Modification. To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$. Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
@@
4.3 Shifts-aware Model-based Offline RL